OpenStudy (anonymous):

Applications of Derivative. An open box with a capacity of 36000 cubic inches is needed. IF the box must be twice as long as it is wide, what dimensions would require the least amount of material?

OpenStudy (dumbcow):

Volume= 36,000 Let x be width, y is height, 2x is length Goal is to minimize surface area SA = (x*2x) + 2(x*y) +2(2x*y) SA = 2x^2 + 6xy ----------------------- Volume = x*2x*y = 2x^2y 2x^2y = 36,000 y = 18,000/x^2 ---------------------------- Substitute into SA equation SA = 2x^2 + 6x(18,000/x^2) SA = 2x^2 + 108,000/x Differentiate and set equal to 0 dA/dx = 4x -108,000/x^2 = 0 4x = 108,000/x^2 4x^3 = 108,000 x^3 = 27,000 x = 30 ------------------------- Solve for y y = 18,000/30^2 y = 20 ------------------------ Dimensions which minimize surface area are width = 30 length = 60 height = 20